OpenStudy (idealist10):

Find an integrating factor for y*dx-x*dy=0 and solve the equation.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Whoops, sorry, forgot about that negative... \[y-x\frac{dy}{dx}=0\] Partial derivatives: \[\frac{\partial}{\partial y}y=1\\ \frac{\partial }{\partial x}x=-1\] Integrating factor: \[\log\mu(x)=\int\frac{1-(-1)}{-x}~dx=-2\int\frac{dx}{x}~~\implies~~\mu(x)=e^{-2\ln x}=\cdots\]

OpenStudy (idealist10):

When you take the partial derivative, isn't it d/dx(-x)=-1 instead of d/dx(x)=-1?

OpenStudy (anonymous):

Yes, sorry. Minor editing error :P

OpenStudy (idealist10):

Thank you so much! I apologize for the slowness of my computer. Please be excused.

OpenStudy (anonymous):

Don't worry about it, and you're very welcome!